B(t) is a brownian motion.

Find the distribution of B(t)=x | B(t+s) = x+k
Answer #1

Now, if X and Y are jointly normal random variables with parameters mu_x, sigma_x^2, mu_y, sigma_y^2, ho , then X | Y is normally distributed with

E[X | Y = y] = mu_x + ho sigma_x (y - mu_y)/sigma_y

Var[X | Y = y] = (1 - ho^2) sigma_x^2

Let X = B(t) and Y = B(t + s), so we have X ~ N(0, t) and Y ~ N(0, t + s) and

Cou(x,y) min(t.s+1) Vtyst


E[X | Y = x+k] = 0 + rac{sqrt{t}}{sqrt{s+t}}. sqrt{t} (x+k - 0)/sqrt{s+t} = rac{t(x+k)}{s+t}

st s+ t

Thus, the distribution of B(t)=x | B(t+s) = x+k is,

X = x | Y = x+k ~ N[ mu = t(x+k)/(s+t) , sigma^2 = st / (s+t)]